ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 127, 2002

Auslander, J. and E. Glasner The distal order of a minimal flow  61

Baracco, L. and G. Zampieri The Boggess–Polking extension theorem for CR functions on manifolds with corners

19

Beidar, K. I., Y. Fong, W.-F. Ke and S. K. Jain An example of a right q-ring

303

Biran, P. and K. Cieliebak Lagrangian embeddings into subcritical Stein manifolds

221

Cieliebak, K. See P. Biran

Farkas, D. R. and L. W. Small Algebras which are nearly finite dimensional and their identities

245

Fong, Y. See K. I. Beidar

Gamburd, A. On the spectral gap for infinite index “congruence” subgroups of SL_{2}(Z)

157

Glasner, E. See J. Auslander

Jain, S. K. See K. I. Beidar

Ke, W.-F. See K. I. Beidar

Koenigsmann, J. Relatively projective groups as absolute Galois groups

93

Koschorke, U. Antipodal vector bundle monomorphisms  29
Man, S. H. and P. F. Smith On chains of prime modules  131

Pestov, V. Ramsey–Milman phenomenon, Urysohn metric spaces, and extremely amenable groups

317

Prado, R. G. A. and F. D. Tall Characterizing ω_{1} and the long line by their topological elementary reflections

81

Ruette, S. Mixing C^{r} maps of the interval without maximal measure 253

Schneider, P. and J. Teitelbaum Banach space representations and Iwasawa theory

359

Small, L. W. See D. R. Farkas

Smith, P. F. See S. H. Man

Stade, E. Archimedean L-factors on GL(n)×GL(n) and generalized Barnes integrals

201

Takahashi, S. Dimension spectra of self-affine sets  1

Takens, F. and E. Verbitskiy Rényi entropies of aperiodic dynamical systems

279

Tall, F. D. See R. G. A. Prado Teitelbaum, J. See P. Schneider Verbitskiy, E. See F. Takens Zampieri, G. See L. Baracco